OpenStudy (anonymous):

Two fair dice are thrown simultaneously. The random variable H is the highest common factor of the two scores. Tabulate the probability ditribution of H, combining together all the possible ways of obtaining the same value.

OpenStudy (anonymous):

The answers are 1= 23/36 2=7/36 3=3/36 4=1/36 5=1/36 6=1/36 and it's probability distributions...

OpenStudy (anonymous):

I'm not sure how they got those answers.

OpenStudy (anonymous):

yeah the problem is weird

OpenStudy (anonymous):

i thought they meant the highest of the 2 dice

OpenStudy (anonymous):

they want the GCF of the two dice

OpenStudy (anonymous):

Yes, they do. I'm not sure how to approach the question with the GCF

OpenStudy (anonymous):

in order for the gcf to be 6, both dice have to be 6 -> 1/6 * 1/6 = 1/36

OpenStudy (anonymous):

similar for 4 and 5

OpenStudy (anonymous):

for the gcf to be 3, both dice must be 3 or 6, but not both 6 because then the gcf would eb 6

OpenStudy (anonymous):

2/6*2/6 - 1/36 = 3/36

OpenStudy (anonymous):

for the gcf to be 2, both dice must be 2,4 or 6, but not both 4 and not both 6

OpenStudy (anonymous):

= 3/6*3/6 - 1/36 - 1/36 = 7/36

OpenStudy (anonymous):

and the rest would have GCF of 1

OpenStudy (anonymous):

Ok, it makes more sense now. Thanks :)